OpenStudy (anonymous):

In professional baseball, a commonly used statistic is winning percentage. To calculate a team's winning percentage, the number of wins is divided by the total games played. Although called a percentage, the winning percentage is usually displayed in decimal form, rounded to three decimal places, like this, for example: .244. The Tacoma Tigers, a new MLB team, were in last place in their division with a record of 29 wins and 43 losses. In other words, their current winning percentage was .

OpenStudy (anonymous):

We are going to find out how many games in a row the Tigers have to win to increase their winning percentage to .500. To set up an equation, consider this: The number of games won in a row would have to be added to both the number of wins and the total number of games, and the result would have to be .500.

OpenStudy (anonymous):

Which of the following equations represents this situation? A. 29+x/72+x= .5 B. 29+x/43+x= .5 C. 29+x/72+x= .403 D. 29+x/43+x= .403

OpenStudy (anonymous):

which one do you think it is?

OpenStudy (gudden):

can u do it now

OpenStudy (anonymous):

I think it is A but Im not sure. :(

OpenStudy (anonymous):

I do to I want to say the answer is A

OpenStudy (anonymous):

Okay well I will go with that. Thanks Patrick!!! I have more questions I am going to post that go along with this kindaaaa

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

Solve the equation (29+x/72+x= .5) to determine how many games the Tigers need to win in a row.

OpenStudy (anonymous):

Ugh can you un-close a question :(